Minutes of the DRAGON meeting of 12 March 2013


Minutes of the DRAGON meeting of 12 March 2013

Present: CR (chair), CA, DH (recorder), PM, GC, BA, LB, BD, AR, JF

1. Hardware

A Conditioning (JF)

At the end of last week the ED1 HV was at about 230 kV.  It was turned down to ~225 kV for the weekend, but on Monday there was a jump of about 5 uA in current drawn.   A similar effect was seen last year.  Then, the SF6 was changed and after some while the voltage slowly recovered to pre-glitch levels for a given anode current.  Now, the current is ~50 uA with the HV only 200 kV instead of 230 kV.   In the accompanying figure: light green = cathode current; tan = anode current; red = set voltage; darker green = readback voltage; light blue = vaccum pressure; dark blue = x-ray counts.

The plan is to continue conditioning for at least one more day to see if there is continuing improvement, as happened last year ("act of faith").   Otherwise a more active intervention will be needed ("auto-da-fe"?)

B. Data acquisition  (GC)

There was a problem with channel 15 of the Wiener HV unit for the BGO detectors.  For now, have just switched to a spare channel.  Was this the problem with BGO #15 last year?  Recollection (JF) is that both HV channels and cables were swapped in testing and the fault stayed with the BGO detector.

The leading edges of analog BGO signals were arriving at the new ADCs less than the 15 ns after the ADC gate (as recommended by in the manual).  The effect on peaks from the calibration source was observed for addition of delays of 16 ns or 64 ns.  Slight narrowing of the peaks was seen with added delay, so 64 ns delays have been added on the analog-in side.  Details in the MIDAS elog of 14:01 on March 8 at http://smaug.triumf.ca:8081/EL/130308.0 .

Pedestals were measured, both with disconnected inputs (pedestals appearing typically around channel 400 (out of 4096)) and with input cables connected but readout triggered by a pulser (typically channel 200).  See the MIDAS elog of  13:39 on March 11 at http://smaug.triumf.ca:8081/EL/.

The BGOs will be re-cabled to make the cable numbering system match the Rev 2 numbering system in analysis codes (i.e. 0,1,2... instead of 1,2,3...).

The problem of the very rare case when ADC data are lost appears to be due to a mistaken belief by the module that it has received a front-panel Reset signal.  

An e-mail about this has been sent to the Caen rep.


2. 7Be(p,p)    (GC)

Energies of elastically scattered protons were calculated as a function of lab angle for incident beam energies of 1 MeV/u and 1.5 MeV/u.  A Si detector 100um thick would (just) stop forward-angle protons at 1 MeV/u.  For higher energies thicker detectors would be needed at forward angles.  Canberra quoted on 100 um detectors at $550 and 300 um detectors at $795.  Suggestion is to use a hybrid of 300 um detectors up to 35 degrees and 100 um detectors at larger angles.


2. 18F(p,gamma) paper  (CA,CR)

Comments have been received from 2 referees, one OK'd the paper for publication, the other wanted a number of changes.   Minor changes have already been made in response, but about a week of work will be required to respond to questions about direct capture and interference of tails of resonances.  The revised version should be circulated to all authors next week. Attached figure shows effect of different interference terms between 3/2+ resonances in 18F(p,g)19Ne. 


3. AOB

-- a draft of the 33S(p,gamma) paper should be circulated next week

-- enriched samples of 80Kr and 34S arrived yesterday


Meeting ended at 15:22